# Diversity of low-mass planet atmospheres in the C-H-O-N-S-Cl system with interior dissolution, non-ideality, and condensation: Application to TRAPPIST-1e and sub-Neptunes

### v2

@@ 3.1.3 Cooled atmosphere and elemental surface inventory
The magma ocean stage establishes the initial volatile inventories in the atmosphere (Section 3.1.2), and as the atmosphere cools and condenses it distributes materials at the planetary surface that can participate in geochemical cycles and fuel a potential biosphere. To this end, we begin with the calculated atmospheric elemental abundances for each of the fully molten mantle cases presented in Section 3.1.2 and cool the atmospheres, isochemically (i.e., assuming a constant bulk atmosphere composition inherited from the magma ocean stage), to 280 K. This is a few kelvin above the freezing point of water at 1 bar and lies between the estimates of the equilibrium temperatures of TRAPPIST-1d and TRAPPIST-1e assuming a null albedo (Delrez et al., 2018). We include the contribution of any graphite present in the abundance of carbon of the high temperature atmosphere, given that graphite remains close to the planetary surface (Keppler & Golabek, 2019) and in equilibrium with the atmosphere. In addition to the aforementioned gas species and graphite (Section 3.1.1), we further allow for the presence of liquid water (H2O), rhombic sulfur (also known as -sulfur), and ammonium chloride (ClH4N) depending on thermodynamic stability, since these condensed phases can be major carriers of elements in the system. The same analysis is applied to the 10% molten mantle cases in Appendix B.
@@
Figure 5 shows the atmospheres at 280 K that were previously in equilibrium with a fully molten mantle and summary statistics are presented in Table 3. The presence of condensates buffers the atmospheric composition and produces well-defined linear trends and clusters in the data. Therefore, and unlike at high temperatures, buffering of gas species fugacities by condensates means that atmospheres at 280 K can be grouped into discontinuous ‘families’, of which we identify four types (see plots of log10SO2 VMR in Figure 5). Among these families, CO2-rich atmospheres are dominant (83% of the total) as revealed by the sharp peak in the frequency distribution at g/mol. These develop from cooling of high temperature atmospheres produced across most of the initial O2 range, except in the most reduced (IW=-3) and oxidised (IW=+3) cases. Initially reduced atmospheres result in a subsidiary peak associated with CH4(g)-rich types (5% of the total) and a tail towards H2(g)-rich atmospheres (1.5% of the total). Another abundant gas in these atmospheres is NH3(g). In general, however, there is a notable dearth of atmospheres with molar masses below about 40 g/mol due to the 1) high C/H starting state caused by H2O(g) dissolution in the magma ocean, and 2) the stability of CO2(g) relative to CO(g) at low temperatures. Finally, oxidised starting conditions (i.e., IW=+3 to +4, see Section 3.1.2) extend the tail of the distribution to high molar masses, defining the fourth family associated with CO2–SO2-bearing atmospheres, which represent 8% of all cases.
These atmospheres can also produce up to 0.3 Earth-ocean equivalents of -sulfur (Figure 5). Crucially, the starting conditions that favour large amounts of sulfur precipitation (oxidised) are different from the conditions that favour significant water, graphite, and ammonium chloride precipitation (IW and below). This arises because S is more abundant in high-temperature atmospheres at oxidising conditions, while C and particularly H are more abundant at more reducing conditions (see Section 3.1.2, Figure 4). There is also a trade-off between graphite and water, since high C/H in the starting conditions promotes larger graphite inventories whereas low C/H forms larger water oceans. In the family of atmospheres with the lowest molar masses () produced by initially low C/H ratios and high H mass fractions, graphite saturation is not reached. The median total atmospheric pressure is roughly 65 bar and the maximum is a few kbar for oxidised starting conditions with high C/H and high initial number of water oceans.
@@ 3.2.1 Context and parameters
Many sub-Neptunes likely harbour massive magma oceans beneath large insulating atmospheres, a configuration that corresponds to so-called ‘gas dwarfs’. Such a structure has been proposed for the canonical sub-Neptune K2-18b (Shorttle et al., 2024; Wogan et al., 2024), although an alternative hypothesis suggests it might support clement conditions at the base of its atmosphere (Madhusudhan et al., 2023). Other sub-Neptune-sized exoplanets are currently being observed with JWST (e.g., TOI-270 d, GJ 1214 b), and preliminary observations reveal absorption features of various gas species . However, whether or not these planets possess a magma ocean surface, a temperate surface that could support liquid water, or no defined surface at all (e.g., Young et al., 2024b), remains unknown. Distinguishing between these scenarios requires self-consistent modelling frameworks that adequately capture the governing physics and chemistry of both the interior and the atmosphere. To this end, Atmodeller allows testing the gas dwarf model with an important addition that has often been neglected in previous work on sub-Neptunes (Schlichting & Young, 2022; Misener et al., 2023; Young et al., 2023; Shorttle et al., 2024; Rogers et al., 2024; Rigby et al., 2024), though see Tian & Heng (2024); Seo et al. (2024): the role of non-ideality, in which the volatile species at the conditions of the magma ocean-atmosphere interface do not obey the ideal gas law. Rather, real gases decouple their partial pressure (which is related to molar abundance, mass balance, and mechanical pressure) from their thermodynamic pressure which regulates equilibrium chemistry and solubility (Section 2.3).
@@
We simulate K2-18 b as a gas dwarf and demonstrate the effects of non-ideality on sub-Neptune atmospheres. K2-18b has a total mass of kg (8.63 M⊕) and an estimated mantle mass of kg (assuming the same core/mantle mass ratio as the Earth) (Benneke et al., 2019). K2-18b’s surface radius, assuming it has an optically thick atmosphere overlying a magma ocean surface, is estimated to be 11225 km (1.76 R⊕; assuming an Earth-like interior composition and the super-Earth mass–radius relation taken from Table 2 of Hakim et al., 2018). The surface temperature at the magma ocean–atmosphere interface is set at 3000 K, We further note that this temperature is at or above the upper limit of the calibration for current solubility laws and real gas EOS. The mantle melt fraction is set to 100%. We consider 6 gas species in the atmosphere, each with their corresponding solubility laws, necessarily assuming the mantle composition is basaltic to remain internally consistent: H2O (Dixon et al., 1995), H2 (Hirschmann et al., 2012), CO (Yoshioka et al., 2019), CO2 (Dixon et al., 1995), and CH4 (Ardia et al., 2013). O2 is also included as a gas species. For the cases with real gases, the Compensated Redlich-Kwong (CORK) model is used for H2O, an ab-initio EOS model for H2 (Chabrier & Debras, 2021), and a corresponding states model for CO, CO2, and CH4 (Shi & Saxena, 1992) (Table 2). O2 is assumed to behave as an ideal gas in all cases, which does not introduce any uncertainty owing to its vanishingly low fugacities under the range of conditions studied.
@@
To explore the effects of non-ideality, we perform two series of simulations in which volatiles can partition between the silicate mantle and the atmosphere. The total hydrogen inventory is varied from 0.1 to 3% of K2-18b’s mass (Figures 6 and 7) assuming both real and ideal gas behaviour. In these simulations, all other parameters are held constant: K, C/H (by mass) = 0.32 ( the solar ratio, as recently inferred for K2-18b, Wogan et al. (2024)) and O2 at IW=-3. The equivalent cases, in which each of the other two variables (IW and C/H ratio) are varied while the others are held constant, are shown in Figures G1 and G2, respectively. All simulation data are available to download (Bower, 2025).
@@ 3.2.2 Volatile storage in the mantle
The key result is that for planets with large volatile inventories, the mass fraction of volatile elements (H+C) retained in the envelope (up to 4% of the planet mass initially added, ) decreases as solubility increases, and decreases even further when real gas behaviour is taken into account. In Figure 6a, a case without solubility (not shown) would produce a 1:1 relationship between the envelope volatile mass fraction (y-axis) and the total volatile mass fraction (x-axis). When ideal gas behaviour and solubility are considered (grey dashed line), the relationship is approximately 2:3, and when both real gas behaviour and solubility are included (grey solid line), the relationship is closer to 1:3. The ideal with solubility case overestimates the volatile mass fraction in the envelope (Figure 6a), and hence the total atmospheric pressure by a factor of 2 compared to the real with solubility case (Figure 7a). Relative to a case without solubility, the envelope volatile mass fraction is reduced by 2/3 when considering both H- and C-bearing species solubilities and real gases. For cases that do consider solubility, assuming a total hydrogen mass equivalent to 3% of K2-18b’s mass results in a 50% decrease of the total atmosphere pressure for the real case (12 GPa) compared to the ideal case (27 GPa). Therefore, these cases demonstrate that both solubility and non-ideality have comparable influences in determining the first-order characteristics of sub-Neptune atmospheres.
The mass fraction of H+C in the envelope predicted by the ideal and real cases with solubility begins to diverge above 0.2–0.3 total volatile mass fraction (compare grey dashed and solid lines, Figure 6). This is because, as the hydrogen mass fraction of the planet increases, the atmospheric surface pressure increases, causing the gas fugacity coefficients, , to deviate more strongly from ideality (Figure 7d). Because the temperatures and pressures far exceed the critical points of the gas species considered here, repulsive forces between the molecules predominate, and volumes (), and hence exceed those of ideal gases. For a constant mass budget of each element, this property increases the fugacity, and hence solubility in silicate liquids, since solubility is proportional to thermodynamic pressure (i.e., fugacity) not partial pressure (Figure 7c). Consequently, the enhanced solubility of volatile elements, primarily hydrogen species (H2 and H2O), limits the amount of H and O (and to a lesser extent C) in the atmosphere and acts to buffer increases in the total atmospheric pressure (Figure 7c,d).
@@
Therefore, CH4 is proportional to whereas scales with . Based on this reasoning, at constant bulk composition, temperature and O2, CH4(g) should become dominant as increases. However, this is not observed in the ideal case, where H2 exceeds CH4 across the entire pressure range (Figure 7a). In ideal cases, H2 remains stable because the addition of gas with C/H (by mass) = 0.32 gives rise to C/H ratios in the atmosphere that are still too low (i.e., 1 by mass; Figure 7b) to form significant quantities of methane (Figure 7a). On the other hand, for real cases, the increase in H2 results in its increased dissolution relative to the ideal case (Figure 7c) while CH4 never dissolves to any substantial extent, causing a commensurate increase in C/H of the atmosphere ( 2.5 by mass, Figure 7b) that is closer to the stoichiometry of CH4 (C/H = 3 by mass). Although CH4 is in the range 103 and H2 20–30, since H2 depends upon H and CH4, the effect of relative differences in essentially cancel one another out.
@@
The aforementioned calculations were performed for a fixed planetary mass and radius. However, these quantities are only determined within limits for sub-Neptunes (at best 15% relative for mass and 3.5% for radius, Luque et al., 2022). Hence we now determine, for a fixed hydrogen mass fraction, C/H, O2 and surface temperature, how variations in a planet’s surface radius or mass can influence the atmospheric speciation we infer when employing real gas EOS. As shown in Figure 8a, in the real case, for a fixed mass of 8.63 M⊕, the atmosphere will be dominated by CH4 below 2 R⊕ and by H2 above 2 R⊕. This contrasts with the ideal scenario, in which H2 is always the dominant species. Similarly, at constant radius (1.76 R⊕), the atmosphere is H2-dominated below 7 M⊕ and CH4-dominated above 7 M⊕ for real gases (Figure 8b). Hence typical uncertainties in mass and radius (Figure 8, grey horizontal bars) can bracket the transition regime between H2 and CH4 dominated atmospheres, highlighting the importance of improving mass and radius measurement uncertainties to accurately interpret atmospheric observations. Mass and radius both influence the atmospheric pressure at the surface, which in turn impact fugacity coefficients, through the relationship:
@@
where is the total mass of the atmosphere, is the acceleration due to gravity, and is the surface radius. Because pressure is proportional to 1/, even small changes in surface radius have tangible effects on total pressure, for a constant atmospheric mass budget. Conversely, the linear dependence of on (and hence on the planetary mass, ) results in a more modest mutual dependence. , and increasing pressure for a constant total mass budget results in higher dissolution (since , where is any given gas species), giving rise to the monotonic decrease in H2 (and hence total ) with increasing or with decreasing in the ideal case. Conversely, increasing leads to increases in (analogous to the mechanism discussed with respect to Equation 6, above), promoting additional H dissolution as molecular H2. The net result is sub-Neptunes are expected to be CH4-rich for cases in which mean densities are roughly 6500 kg/m3 at IW=-3, H=3% , and bulk C/H=100 solar. Figures G1 and G2 illustrate the effects of varying oxygen fugacity from (at constant C/H = 0.32 and H mass fraction = 1% ) and total C/H mass ratio (at constant O2 = IW=-3 and H mass fraction = 1% ), respectively, on volatile speciation in the atmosphere and dissolution in the interior assuming real and ideal gas behaviour, for K2-18b-like cases.

### v3

@@ 3.1.3 Cooled atmosphere and elemental surface inventory
The magma ocean stage establishes the initial volatile inventories in the atmosphere (Section 3.1.2), and as the atmosphere cools and condenses it distributes materials at the planetary surface that can participate in geochemical cycles and fuel a potential biosphere. To this end, we begin with the calculated atmospheric elemental abundances for each of the fully molten mantle cases presented in Section 3.1.2 and cool the atmospheres, isochemically (i.e., assuming a constant bulk atmosphere composition inherited from the magma ocean stage), to 280 K. This is a few kelvin above the freezing point of water at 1 bar and lies between the estimates of the equilibrium temperatures of TRAPPIST-1d and TRAPPIST-1e assuming a null albedo (Delrez et al., 2018). We include the contribution of any graphite present in the abundance of carbon of the high-temperature atmosphere, given that graphite remains close to the planetary surface (Keppler & Golabek, 2019) and in equilibrium with the atmosphere. In addition to the aforementioned gas species and graphite (Section 3.1.1), we further allow for the presence of liquid water (H2O), rhombic sulfur (also known as -sulfur), and ammonium chloride (ClH4N) depending on thermodynamic stability, since these condensed phases can be major carriers of elements in the system. The same analysis is applied to the 10% molten mantle cases in Appendix B.
@@
Figure 5 shows the atmospheres at 280 K that were previously in equilibrium with a fully molten mantle and summary statistics are presented in Table 3. The presence of condensates buffers the atmospheric composition and produces well-defined linear trends and clusters in the data. Therefore, and unlike at high temperatures, buffering of gas species fugacities by condensates means that atmospheres at 280 K can be grouped into discontinuous ”families,” of which we identify four types (see plots of log10SO2 VMR in Figure 5). Among these families, CO2-rich atmospheres are dominant (83% of the total) as revealed by the sharp peak in the frequency distribution at g mol-1. These develop from cooling of high temperature atmospheres produced across most of the initial O2 range, except in the most reduced (IW = -3) and oxidized (IW = +3) cases. Initially reduced atmospheres result in a subsidiary peak associated with CH4(g)-rich types (5% of the total) and a tail toward H2(g)-rich atmospheres (1.5% of the total). Another abundant gas in these atmospheres is NH3(g). In general, however, there is a notable dearth of atmospheres with molar masses below about 40 g mol-1 due to the (1) high C/H starting state caused by H2O(g) dissolution in the magma ocean, and (2) the stability of CO2(g) relative to CO(g) at low temperatures. Finally, oxidized starting conditions (i.e., IW = +3 to +4, see Section 3.1.2) extend the tail of the distribution to high molar masses, defining the fourth family associated with CO2–SO2-bearing atmospheres, which represent 8% of all cases.
These atmospheres can also produce up to 0.3 Earth-ocean equivalents of -sulfur (Figure 5). Crucially, the starting conditions that favor large amounts of sulfur precipitation (oxidized) are different from the conditions that favor significant water, graphite, and ammonium chloride precipitation (IW and below). This arises because S is more abundant in high-temperature atmospheres at oxidizing conditions, while C and particularly H are more abundant at more reducing conditions (see Section 3.1.2, Figure 4). There is also a trade-off between graphite and water, since high C/H in the starting conditions promotes larger graphite inventories, whereas low C/H forms larger water oceans. In the family of atmospheres with the lowest molar masses () produced by initially low C/H ratios and high H mass fractions, graphite saturation is not reached. The median total atmospheric pressure is roughly 65 bar and the maximum is a few kbar for oxidized starting conditions with high C/H and high initial number of water oceans.
@@ 3.2.1 Context and parameters
Many sub-Neptunes likely harbor massive magma oceans beneath large insulating atmospheres, a configuration that corresponds to so-called ”gas dwarfs.” Such a structure has been proposed for the canonical sub-Neptune K2-18b (Shorttle et al., 2024; Wogan et al., 2024), although an alternative hypothesis suggests it might support clement conditions at the base of its atmosphere (Madhusudhan et al., 2023). Other sub-Neptune-sized exoplanets are currently being observed with JWST (e.g., TOI-270 d, GJ 1214 b), and preliminary observations reveal absorption features of various gas species . However, whether or not these planets possess a magma ocean surface, a temperate surface that could support liquid water, or no defined surface at all (e.g., Young et al., 2024b), remains unknown. Distinguishing between these scenarios requires self-consistent modeling frameworks that adequately capture the governing physics and chemistry of both the interior and the atmosphere. To this end, Atmodeller allows us to test the gas dwarf model with an important addition that has often been neglected in previous work on sub-Neptunes (Schlichting & Young, 2022; Misener et al., 2023; Young et al., 2023; Shorttle et al., 2024; Rogers et al., 2024; Rigby et al., 2024), though see Tian & Heng (2024); Seo et al. (2024): the role of nonideality, in which the volatile species at the conditions of the magma ocean–atmosphere interface do not obey the ideal gas law. Rather, real gases decouple their partial pressure (which is related to molar abundance, mass balance, and mechanical pressure) from their thermodynamic pressure, which regulates equilibrium chemistry and solubility (Section 2.3).
@@
We simulate K2-18 b as a gas dwarf and demonstrate the effects of nonideality on sub-Neptune atmospheres. K2-18b has a total mass of kg (8.63 ) and an estimated mantle mass of kg (assuming the same core/mantle mass ratio as the Earth) (Benneke et al., 2019). K2-18b’s surface radius, assuming it has an optically thick atmosphere overlying a magma ocean surface, is estimated to be 11225 km (1.76 ; assuming an Earth-like interior composition and the super-Earth mass–radius relation taken from Table 2 of Hakim et al., 2018). The surface temperature at the magma ocean–atmosphere interface is set at 3000 K, We further note that this temperature is at or above the upper limit of the calibration for current solubility laws and real gas EOS. The mantle melt fraction is set to 100%. We consider six gas species in the atmosphere, each with their corresponding solubility laws, necessarily assuming the mantle composition is basaltic to remain internally consistent: H2O (Dixon et al., 1995), H2 (Hirschmann et al., 2012), CO (Yoshioka et al., 2019), CO2 (Dixon et al., 1995), and CH4 (Ardia et al., 2013). O2 is also included as a gas species. For the cases with real gases, the CORK model is used for H2O, an ab initio EOS model for H2 (Chabrier & Debras, 2021), and a corresponding states model for CO, CO2, and CH4 (Shi & Saxena, 1992) (Table 2). O2 is assumed to behave as an ideal gas in all cases, which does not introduce any uncertainty owing to its vanishingly low fugacities under the range of conditions studied.
@@
To explore the effects of nonideality, we perform two series of simulations in which volatiles can partition between the silicate mantle and the atmosphere. The total hydrogen inventory is varied from 0.1% to 3% of K2-18b’s mass (Figures 6 and 7) assuming both real and ideal gas behavior. In these simulations, all other parameters are held constant: K, C/H (by mass) = 0.32 ( the solar ratio, as recently inferred for K2-18b, Wogan et al. (2024)) and O2 at IW = -3. The equivalent cases, in which each of the other two variables (IW and C/H ratio) are varied while the others are held constant, are shown in Figures G1 and G2, respectively. All simulation data are available to download (Bower, 2025).
@@ 3.2.2 Volatile storage in the mantle
The key result is that for planets with large volatile inventories, the mass fraction of volatile elements (H+C) retained in the envelope (up to 4% of the planet mass initially added, ) decreases as solubility increases, and decreases even further when real gas behavior is taken into account. In Figure 6a, a case without solubility (not shown) would produce a 1:1 relationship between the envelope volatile mass fraction (-axis) and the total volatile mass fraction (-axis). When ideal gas behavior and solubility are considered (gray dashed line), the relationship is approximately 2:3, and when both real gas behavior and solubility are included (gray solid line), the relationship is closer to 1:3. The ideal with solubility case overestimates the volatile mass fraction in the envelope (Figure 6a), and hence the total atmospheric pressure by a factor of 2 compared to the real with solubility case (Figure 7a). Relative to a case without solubility, the envelope volatile mass fraction is reduced by 2/3 when considering both H- and C-bearing species solubilities and real gases. For cases that do consider solubility, assuming a total hydrogen mass equivalent to 3% of K2-18b’s mass results in a 50% decrease of the total atmosphere pressure for the real case (12 GPa) compared to the ideal case (27 GPa). Therefore, these cases demonstrate that both solubility and nonideality have comparable influences in determining the first-order characteristics of sub-Neptune atmospheres.
The mass fraction of H+C in the envelope predicted by the ideal and real cases with solubility begins to diverge above 0.2–0.3 total volatile mass fraction (compare gray dashed and solid lines, Figure 6). This is because, as the hydrogen mass fraction of the planet increases, the atmospheric surface pressure increases, causing the gas fugacity coefficients, , to deviate more strongly from ideality (Figure 7d). Because the temperatures and pressures far exceed the critical points of the gas species considered here, repulsive forces between the molecules predominate, and volumes (), and hence exceed those of ideal gases. For a constant mass budget of each element, this property increases the fugacity, and hence solubility in silicate liquids, since solubility is proportional to thermodynamic pressure (i.e., fugacity) not partial pressure (Figure 7c). Consequently, the enhanced solubility of volatile elements, primarily hydrogen species (H2 and H2O), limits the amount of H and O (and to a lesser extent C) in the atmosphere and acts to buffer increases in the total atmospheric pressure (Figure 7c,d).
@@
Therefore, CH4 is proportional to , whereas scales with . Based on this reasoning, at constant bulk composition, temperature and O2, CH4(g) should become dominant as increases. However, this is not observed in the ideal case, where H2 exceeds CH4 across the entire pressure range (Figure 7a). In ideal cases, H2 remains stable because the addition of gas with C/H (by mass) = 0.32 gives rise to C/H ratios in the atmosphere that are still too low (i.e., 1 by mass; Figure 7b) to form significant quantities of methane (Figure 7a). On the other hand, for real cases, the increase in H2 results in its increased dissolution relative to the ideal case (Figure 7c) while CH4 never dissolves to any substantial extent, causing a commensurate increase in C/H of the atmosphere ( 2.5 by mass, Figure 7b) that is closer to the stoichiometry of CH4 (C/H = 3 by mass). Although CH4 is in the range 103 and H2 20–30, since H2 depends upon H and CH4, the effect of relative differences in essentially cancel one another out.
@@
The aforementioned calculations were performed for a fixed planetary mass and radius. However, these quantities are only determined within limits for sub-Neptunes (at best 15% relative for mass and 3.5% for radius, Luque et al., 2022). Hence, we now determine, for a fixed hydrogen mass fraction, C/H, O2 and surface temperature, how variations in a planet’s surface radius or mass can influence the atmospheric speciation we infer when employing real gas EOS. As shown in Figure 8a, in the real case, for a fixed mass of 8.63 , the atmosphere will be dominated by CH4 below 2 and by H2 above 2 . This contrasts with the ideal scenario, in which H2 is always the dominant species. Similarly, at constant radius (1.76 R⊕), the atmosphere is H2 dominated below 7 M⊕ and CH4 dominated above 7 M⊕ for real gases (Figure 8b). Hence, typical uncertainties in mass and radius (Figure 8, gray horizontal bars) can bracket the transition regime between H2 and CH4 dominated atmospheres, highlighting the importance of improving mass and radius measurement uncertainties to accurately interpret atmospheric observations. Mass and radius both influence the atmospheric pressure at the surface, which in turn impact fugacity coefficients, through the relationship:
@@
where is the total mass of the atmosphere, is the acceleration due to gravity, and is the surface radius. Because pressure is proportional to 1/, even small changes in surface radius have tangible effects on total pressure, for a constant atmospheric mass budget. Conversely, the linear dependence of on (and hence on the planetary mass, ) results in a more modest mutual dependence. , and increasing pressure for a constant total mass budget results in higher dissolution (since , where is any given gas species), giving rise to the monotonic decrease in H2 (and hence total ) with increasing or with decreasing in the ideal case. Conversely, increasing leads to increases in (analogous to the mechanism discussed with respect to Equation 6, above), promoting additional H dissolution as molecular H2. The net result is sub-Neptunes are expected to be CH4-rich for cases in which mean densities are roughly 6500 kg m-3 at IW = -3, H=3% , and bulk C/H=100 solar. Figures G1 and G2 illustrate the effects of varying oxygen fugacity from -6 ¡ IW ¡ 0 (at constant C/H = 0.32 and H mass fraction = 1% ) and total C/H mass ratio (at constant O2 = IW = -3 and H mass fraction = 1% ), respectively, on volatile speciation in the atmosphere and dissolution in the interior assuming real and ideal gas behavior, for K2-18b-like cases.